OpenStudy (anonymous):

I really need help in connections Geometry Unit 5 lesson 9 Surface Area and Volume Unit test please?!?!? Im having a really hard time and anyone willing to help it would be greatly appreciated!!

OpenStudy (moonlitfate):

Can you give an example of one of the problems that you're having issues with?

OpenStudy (anonymous):

I'm basically having issues with most of them. Math definitely isn't my stronger subject.

OpenStudy (anonymous):

Find the lateral area and surface area of the cone. Round the answers to the nearest tenth. The figure is not drawn to scale. (1 point) L.A. = 395.8 ft2; S.A. = 791.7 ft2 L.A. = 1,583.4 ft2; S.A. = 1,244.1 ft2 L.A. = 622.0 ft2; S.A. = 791.7 ft2 L.A. = 791.7 ft2; S.A. = 1,244.1 ft2

OpenStudy (anonymous):

That's only one of them.

OpenStudy (moonlitfate):

Okay, well' let's do that one first. :)

OpenStudy (anonymous):

Please. xD

OpenStudy (moonlitfate):

-- So, I flipped the cone there. And if you look at the cone, the base is a circle. And if you draw a line from the radius to the top of the cone, you get a triangle. Do you see that? [And sorry if my explanations aren't so great; not so great with Geometry]

OpenStudy (anonymous):

That's alright- me either, obviously! Yes, I see it. Isn't there a formula for calculate the surface and linear areas of a cone though? I was having a hard time recalling it.

OpenStudy (moonlitfate):

Yes. :)

OpenStudy (anonymous):

Do you know what the formulas are?

OpenStudy (moonlitfate):

Surface Area of a cone is:\[SA = \pi r s + \pi r^2\] It's basically the area of the base and the area of the cone added together. r = the radius, and s = the side of the that we were given.

OpenStudy (anonymous):

Okay, so I can try to solve it then.

OpenStudy (anonymous):

r = 12 s = 21 3.14 x 12 x 21 = 791.28

OpenStudy (anonymous):

3.14 x 12^2 = 3.14 x 144 3.14 x 144 = 452.16

OpenStudy (moonlitfate):

Lateral Area of a cone is: \[LA = \pi r L\] where r is the radius, L is the slant height of the cone.

OpenStudy (anonymous):

wouldn't the slant height be 21?

OpenStudy (anonymous):

I believe that the answer is D. Thanks for your help. (:

OpenStudy (moonlitfate):

\[SA = \pi(12)(21)+\pi(12)^2\]\[=791.681 + 452.389\] = 1244.1 And yes, the slant height is 21. :)

OpenStudy (anonymous):

Thanks for your help (:

OpenStudy (moonlitfate):

You're welcome Does it make sense now?